OpenStudy (anonymous):

divide and check (36t^6-27t^5-9t^20/(-9t^2)

OpenStudy (pfenn1):

Is this your equation? \[\frac{36t^6-27t^5-9t^{20}}{-9t^2}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wait -9t^2 not 9t^20

OpenStudy (pfenn1):

ok then this?\[\frac{36t^6-27t^5-9t^{2}}{-9t^2}\]

OpenStudy (pfenn1):

\[\frac{36t^6-27t^5-9t^{2}}{-9t^2}=\frac{36t^6}{-9t^2}+\frac{-27t^5}{-9t^2}+\frac{-9t^2}{-9t^2}\] can you simplify this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes

OpenStudy (pfenn1):

So what do you get for the simplified version?

OpenStudy (anonymous):

4t^3+3t^2+1t^1 then 74 i think

OpenStudy (anonymous):

is this correct

OpenStudy (pfenn1):

It isn't what I got. \[\frac{36t^6}{-9t^2}=-4t^4\] I'm not sure how you got \[4t^3\]

OpenStudy (pfenn1):

\[\frac{-27t^5}{-9t^2}=3t^3\]\[\frac{-9t^2}{-9t^2}=1\]

OpenStudy (pfenn1):

\[-4t^4+3t^3+1\]

OpenStudy (anonymous):

i didn't subtract almost got it thank you very much

OpenStudy (anonymous):

so this is my answer